# SageAttention2 Technical Report: Accurate 4 Bit Attention for Plug-and-play Inference Acceleration

### v1

@@ 4.2 Speed and Accuracy of Kernels
Speed. We conduct experiments to compare the Speed of SageAttention2 against baselines using configurations with headdim=64, headdim=128, and headdim=256, both with and without Causal Mask (Vaswani, 2017). Specifically, Figure 8, Figure 9, and Figure 10 show the speed of SageAttention2 and baselines across varying sequence lengths on RTX4090. These results indicate that SageAttention2 achieves a peak of 485 TOPS and is 3.1x faster than FlashAttention2 and 5.4x faster than xformers. Figure 11, Figure 12, and Figure 13 illustrate the results on L20 GPU, indicating that SageAttention2 achieves a peak of 288 TOPS and is 2.7x faster than FlashAttention2 and 4.6x faster than xformers.
Accuracy. Table 10 and Table 10 show that the average and worst accuracy of different methods with INT4 $Q,K$ and FP8 $P,V$ across all layers of CogvideoX. The results indicate that the accuracy of SageAttn-4b is superior to other baselines.
Figure 15: End-to-end performance using different ratio of SageAttn-8b.
@@ 4.3 End-to-end Performance
Metrics loss. We assessed the end-to-end metrics of various models using SageAttention2 compared to using attention in full precision. Detailed evaluation results are presented in Table 11 for Llama2, Llama3.1, GLM4, CogvideoX, Open-Sora, Flux, and TIMM, respectively. The results indicate that SageAttn-4b outperforms all baselines and maintains most of the end-to-end accuracy across all models. Additionally, using adaptive quantization techniques, SageAttn-mix achieves performance comparable to full-precision attention.
@@ 4.4 Ablation Study
Adaptive quaitzation. To analyze the impact of using different ratio of SageAttn-8b in adaptive quantization, Table 15 shows the changes in the perplexity of Llama3.1 under various ratio of SageAttn-8b. It can be observed that even if SageAttn-4b is used exclusively, the overall end-to-end representation is sufficiently good, and the higher the ratio of SageAttn-8b used, the closer the accuracy approaches that of using full-precision attention.
Overhead of smoothing Q, K. Note that the overhead of smoothing $Q,K$ only contains $Q-\overrightarrow{Q_{m}}$, $K-\mathrm{mean}(K)$, and $\overrightarrow{Q_{m}}K^{\top}$. The speed overhead of them accounts for about 3.5% of the attention kernel.
@@ 5 Conclusion and Future Work
We introduce SageAttention2, an efficient and accurate 4-bit quantization method for attention. First, we propose to quantize $(Q,K)$ to INT4 in a warp-level granularity and quantize $(\widetilde{P},V)$ to FP8. Second, we propose a method to smooth matrixes $Q$ and $V$, enhancing the accuracy of attention with INT4 $QK$ and FP8 $PV$. Third, we analyze the quantization accuracy across timesteps and layers, then propose an adaptive quantization method to ensure the end-to-end metrics over various models. Our implementation is faster than FlashAttention2 and xformers by approximately 3.1x and 5.4x on RTX4090, respectively. Extensive testing confirms that our approach maintains end-to-end metrics across various models, including language, image, and video generation.

### v3

@@ 4.2 Speed and Accuracy of Kernels
Speed. We conduct experiments to compare the Speed of SageAttention2 against baselines using configurations with headdim=64 and headdim=128, both with and without Causal Mask (Vaswani, 2017). Specifically, Figure 8 and Figure 9, show the speed of SageAttention2 and baselines across varying sequence lengths on RTX4090. These results indicate that SageAttn2-4b and SageAttn2-8b are approximately 3x and 2.7x faster than FlashAttention2, and about 4.5x and 4x faster than xformers, respectively.
Accuracy. Table 9 and Table 9 show that the average and worst accuracy of different methods with INT4 $Q,K$ and FP8 $P,V$ across all layers of CogvideoX. The results indicate that the accuracy of SageAttn2-4b is superior to other baselines.
Table 12: End-to-end generation latency using SageAttention2 (The latency of Llama3.1 is the time to first token generation using different sequence lengths).
Model GPU Original SageAttn2-8b SageAttn2-4b
CogvideoX (2B) RTX4090 86 s 54 s 52 s
CogvideoX (1.5-5B) RTX4090 1040 s 577 s 555 s
HunyuanVideo L20 2221 s 1486 s 1435 s
Mochi L20 2336 s 1316 s 1190 s
Llama3.1 (48K tokens) RTX4090 9.2 s 5.7 s 5.6 s
Llama3.1 (100K tokens) L20 39.9 s 25.4 s 23.2 s
@@ 4.3 End-to-end Performance
Metrics loss. We assessed the end-to-end metrics of various models using SageAttention2 compared to using attention in full precision. Detailed evaluation results are presented in Table 10 for Llama2, Llama3.1, GLM4, CogvideoX (2B), CogvideoX (1.5-5B), HunyuanVideo, Mochi, Flux, Stable-Diffusion3.5, and TIMM, respectively. The results indicate that SageAttn2-4b outperforms all baselines and maintains most of the end-to-end accuracy across all models. Additionally, SageAttn2-8b incurs almost no metrics loss across various models.
Visible examples. Figure12 and Figure 13 show some visible comparision examples from HunyuanVideo and CogvideoX (1.5-5B). We can observe that SageAttn2-8b does not introduce any visible differences compared to full-precision attention, whereas SageAttn2-4b has minor differences compared to full-precision attention but is much better than the baselines.
End-to-end speedup. We compared the original generation latency and the latency using SageAttention2 for models with long sequence lengths in Table 12, observing significant speedup effects. For instance, SageAttention2 achieved a 1.8x speedup in CogvideoX (1.5-5B) without any metrics loss (SageAttn2-8b). SageAttn2-4b further accelerated these models but with a little metrics loss.
Smooth V Cos Sim $\uparrow$ Relative L1 $\downarrow$ RMSE $\downarrow$
98.25% 0.1980 0.2387
99.75% 0.0406 0.0773
@@ 4.4 Ablation Study
Overhead of smoothing Q, K. Note that the overhead of smoothing $Q,K$ only contains $Q-\overrightarrow{Q_{m}}$, $K-\mathrm{mean}(K)$, and $\overrightarrow{Q_{m}}K^{\top}$. The speed overhead of smoothing $Q,K$ accounts for about 3.5% of the attention kernel.
Benefit of smoothing V. Table 13 shows the attention accuracy on real tensors sampled from CogvideoX with and without smoothing $V$. It demonstrates that smoothing V could improve the accuracy of SageAttention2 when quantizing $Q,K$ to INT4 and $\widetilde{P},V$ to FP8.
@@ 5 Conclusion and Future Work
We introduce SageAttention2, an efficient and accurate quantization method for attention. First, we propose to quantize matrixes $(Q,K)$ to INT4 in a thread-level granularity and quantize matrixes $(\widetilde{P},V)$ to FP8. Second, we propose a method to smooth $Q$, enhancing the accuracy of INT4 $QK^{\top}$. Third, we propose to use an FP32 Matmul buffer for $PV$ to enhance the accuracy of FP8 $\widetilde{P}V$. Our implementation is faster than FlashAttention2 and xformers by approximately 3x and 5x on RTX4090, respectively. Extensive testing confirms that our approach maintains end-to-end metrics across various models, including language, image, and video generation.